OpenStudy (anonymous):

A hot-air balloon rises 90 feet in its first minute of flight. In each succeeding minute, it rises only 90% as far as it did during the preceeding minute. What is the final height of the balloon when it stops rising?

OpenStudy (amistre64):

90 + 90(.9) + 90(.9)(.9) + 90(.9)(.9)(.9) + ....

OpenStudy (amistre64):

90/(1-.90)

OpenStudy (anonymous):

How far do you keep going?

OpenStudy (amistre64):

infinity :)

OpenStudy (amistre64):

\[90\frac{1-.9^{\infty}}{1-.9}=90\frac{1}{1-.9}\]

OpenStudy (anonymous):

So 900?

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

Thank you! :)

OpenStudy (amistre64):

yw